OpenStudy (anonymous):

simplify2 sqrt 7 times 3 sqrt 14

OpenStudy (anonymous):

\[2\sqrt{7}*3\sqrt{7}*\sqrt{2} \rightarrow 6*7*\sqrt{2} \rightarrow 42\sqrt{2}\]